An 8.00 kg shell at rest explodes in two fragments, with mass ratio of 3:2. If the lighter fragment


gains 90.0 3 of kinetic energy from the explosion, how much kinetic energy does the heavier fragment gain?

1
Expert's answer
2022-04-18T17:31:33-0400

According to the momentum conservation law, the momentum "p" of each fragment is the same. By definition, the kinetic energies of them is given as follows:


"K_1 = \dfrac{p^2}{2m_1}\\
K_2 = \dfrac{p^2}{2m_2}"

where "m_1, m_2" are the masses of the first and second fragment respectively, and "m_2\/m_1 = 3\/2".

Let "K_1 = 90.0J". The ratio of kinetic energies is:


"\dfrac{K_2}{K_1} = \dfrac{p^2}{2m_2} /\dfrac{p^2}{2m_1} = \dfrac{m_1}{m_2}\\
K_2 = K_1\dfrac{m_1}{m_2}\\
K_2 = 90.0J\cdot \dfrac{2}{3} = 60.0J"

Answer. 60.0J.